OpenStudy (anonymous):

1. Using complete sentences, describe how you would analyze the zeros of the polynomial function f(x) = 7x5 + 15x4 – x3 + 4x2 – 6x – 11 using Descartes’ Rule of Signs. Be sure to provide the answer in your explanation.

jimthompson5910 (jim_thompson5910):

how many sign changes are there in f(x)

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so there are at most 3 positive real roots

jimthompson5910 (jim_thompson5910):

now evaluate f(-x) and tell me what you get

OpenStudy (anonymous):

uh i dont know how.

jimthompson5910 (jim_thompson5910):

start with f(x) and replace all copies of x you see with -x then simplify

OpenStudy (anonymous):

f(-x) = -7x5 - 15x4 + x3 - 4x2 + 6x + 11 is this what you meant? Original function: f(x) = 7x5 + 15x4 – x3 + 4x2 – 6x – 11

jimthompson5910 (jim_thompson5910):

f(x) = 7x^5 + 15x^4 – x^3 + 4x^2 – 6x – 11 f(-x) = 7(-x)^5 + 15(-x)^4 – (-x)^3 + 4(-x)^2 – 6(-x) – 11 f(-x) = -7x^5 + 15x^4 + x^3 + 4x^2 + 6x – 11

OpenStudy (anonymous):

so i sorta did it right?

jimthompson5910 (jim_thompson5910):

a bit, but you have signs off in some places

OpenStudy (anonymous):

shouldnt 4x^2 be a negative ?

jimthompson5910 (jim_thompson5910):

nope because (-x)^2 = x^2

jimthompson5910 (jim_thompson5910):

so +4(-x)^2 = +4x^2

OpenStudy (anonymous):

oh okay.

jimthompson5910 (jim_thompson5910):

this is true for any even exponent

OpenStudy (anonymous):

ohkay so like is that it?

jimthompson5910 (jim_thompson5910):

now you count the sign changes in f(-x)

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

so there are at most 2 negative real roots

jimthompson5910 (jim_thompson5910):

what is the degree of this function

OpenStudy (anonymous):

of the original or new one?

jimthompson5910 (jim_thompson5910):

either one, but mainly the original

OpenStudy (anonymous):

5?

jimthompson5910 (jim_thompson5910):

the degree tells you what about the total number of roots

OpenStudy (anonymous):

so 5?

jimthompson5910 (jim_thompson5910):

good, so if you have a max of 3 positive real roots and a max of 2 negative real roots how many complex roots do you have?

OpenStudy (anonymous):

none

jimthompson5910 (jim_thompson5910):

now let's say you only had 2 positive real roots and a max of 2 negative real roots how many complex roots do you have?

OpenStudy (anonymous):

still none right?

jimthompson5910 (jim_thompson5910):

you have 2+2 = 4 roots so far but there are 5 total, so you have room for 1 more

jimthompson5910 (jim_thompson5910):

but you can't have just 1 complex root in the form a+bi since they always come in pairs

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yea.

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

so you could have 5 real roots, with no imaginary roots at all or you could have 3 real roots with 2 imaginary roots or you could have 1 real root with 4 imaginary roots

jimthompson5910 (jim_thompson5910):

and the real roots break down like this 3 max positive roots 2 max negative roots

OpenStudy (anonymous):

okay.